OpenStudy (anonymous):

Find the first, fourth, and eighth terms of the sequence. A(n)=-5 x 2^n-1. HELP PLEASE???????????

OpenStudy (anonymous):

replace \(n\) by 1 for the first term, bet \[A(1)=-5\times 2^{1-1}=-5\times 2^0=-5\] (since \(2^0=1\))

OpenStudy (anonymous):

now replace \(n\) by 4 and get \[A(4)=-5\times 2^{4-1}=-5\times 2^3=-5\times 8=-40\]

OpenStudy (anonymous):

your on your own with the last one, but the procedure is the same as for the first two